### v1

# Problem Solved? Information Extraction Design Space for Layout-Rich Documents using LLMs

Abstract
This paper defines and explores the design space for information extraction (IE) from layout-rich documents using large language models (LLMs). The three core challenges of layout-aware IE with LLMs are 1) data structuring, 2) model engagement, and 3) output refinement. Our study delves into the sub-problems within these core challenges, such as input representation, chunking, prompting, and selection of LLMs and multimodal models. It examines the outcomes of different design choices through a new layout-aware IE test suite, benchmarking against the state-of-art (SoA) model LayoutLMv3. The results show that the configuration from one-factor-at-a-time (OFAT) trial achieves near-optimal results with $14.1$ points F1-score gain from the baseline model, while full factorial exploration yields only a slightly higher $15.1$ points gain at 36$\sim$ 36$\times$ 36 × greater token usage. We demonstrate that well-configured general-purpose LLMs can match the performance of specialized models, providing a cost-effective alternative. Our test-suite is freely available at https://github.com/gayecolakoglu/LayIE-LLM.
Problem Solved? Information Extraction Design Space for Layout-Rich Documents using LLMs
Gaye Colakoglu1,2  Gürkan Solmaz2  Jonathan Fürst1 1Zurich University of Applied Sciences, Switzerland 2NEC Laboratories Europe, Heidelberg, Germany colgay01@students.zhaw.ch, gurkan.solmaz@neclab.eu, jonathan.fuerst@zhaw.ch
1 Introduction
Information extraction (IE) entails extracting structured data, such as names, dates, or financial figures, from unstructured documents. Within organizations, key information often resides in layout-rich documents (LRDs) such as reports and presentations that combine visual elements (e.g., charts, tables) with textual structure and content Park et al. (2019); Wang et al. (2023a); Zmigrod et al. (2024b). LRDs challenge traditional natural language processing (NLP) techniques, which are designed for plain texts Cui et al. (2021); Tang et al. (2023).
Figure 1: Design space for information extraction from layout-rich documents using LLMs. The goal is to extract all relevant information to the target data schema with correct mapping.
Recent layout-aware models at the intersection of NLP and Computer Vision (CV) address this gap by including visual and structural features to improve information extraction from LRDs (e.g., LayoutLMv1-v3 Xu et al. (2020b, a); Huang et al. (2022), GraphDoc Zhang et al. (2022), DocFormer Appalaraju et al. (2021)). However, these models require substantial dataset-specific fine-tuning; users must manually annotate a training dataset with bounding boxes and extraction elements for each new document set.
While LLMs show significant potential, there exist open questions regarding about their adoption for IE from LRDs. In which form should the document content be provided to the LLM? Which methods are most effective for in-context learning (ICL) and instruction tuning? Do larger multimodal LLMs like GPT-4o provide advantages over traditional Optical Character Recognition (OCR) and text-based models? How coherent is the LLM output, and what post-processing is needed? How does the performance of open- and closed-sourced LLMs of different scales compare? How does performance compare to fine-tuned SoA layout-aware models?
This paper addresses the questions above, by systematically exploring the design space for IE with LLMs from LRDs. We implement and evaluate a range of preprocessing, chunking, prompting, and post-processing techniques alongside ICL strategies. Furthermore, we experimentally assess the performance of diverse LLMs and LayoutLMv3 Huang et al. (2022), a fine-tuned layout-aware model.
Our results reveal multiple important insights. First, current, general LLMs can easily compete with SoA fine-tuned models such as LayoutLMv3 with the additional benefit to not require any training data. Second, instead of fine-tuning through data, LLMs require tuning of the IE pipeline to achieve a competitive performance—the gap between a best-practices baseline and the tuned configuration achieved by our lightweight OFAT method is 14.1 points in F1. Third, while purely text-based LLMs achieve a competitive performance through our method, multi-modal LLMs that directly combine textual with visual features still achieve higher performance with the disadvantage of higher costs (token use and API costs) and less transparency.
Shortly, our contributions are as follows:
  • Introducing the Design Space of IE from LRDs using LLMs, consisting of the three core challenges: 1) Data structuring, 2) Model engagement, and 3) Output refinement (Sec. 2).
  • Development of a layout-aware IE test suite for analyzing effects of OCR- and text-based inputs, chunk sizes, few-shots and CoT prompting, LLM model selection, decoding, entity mapping, data cleaning, and F1-score-based benchmarking using exact, substring, and fuzzy matches (Sec. 3).
  • Comprehensive evaluation using GPT-4o, GPT3.5, and LLaMA3 models, GPT4-vision, and the SoA LayoutLMv3 model (Sec. 4).
  • We open-source all our code and experimental results (data accessible in Wang et al. (2023b)). Our test suite can be used and adapted by others to tune their IE pipelines to their datasets and LLMs (https://github.com/gayecolakoglu/LayIE-LLM).
2 Design Space of IE from Layout-rich Documents with LLMs
Task Definition.
IE from LRDs involves identifying and extracting information from documents where textual content is intertwined with complex visual layouts and mapping them into structured information instances such that
$\text{IE}:(D,S)\rightarrow E$ (1)
,where
  • D represents the set of LRDs, each with content and layout information.
  • S is the target schema that defines the set of slots to be filled. Each slot is defined by an attribute (key) $a_{i}$ and its corresponding data type (domain) $T_{i}$, such that $S=\{(a_{1},T_{1}),(a_{2},T_{2}),...,(a_{k},T_{k})\}$.
  • Finally, E represents the set of extracted information instances, where each instance is a set of slot-value pairs derived from a document in D, leveraging both content and layout to determine the correct values for the slots in S. Each value in an instance must conform to the data type $T$, specified in the schema for that attribute.
2.1 Using LLMs for Information Extraction
IE systems that utilize LLMs have to tackle the three main challenge areas that we consider as part of the Design Space: Data Structuring, Model Engagement, and Output Refinement. Each stage, in their respective order, plays an important role for having an IE system with satisfactory accuracy and robustness.
Data Structuring. For multimodal LLMs, LRDs can be directly given as input. On the other hand, for purely text-based LLM, the input documents must be transformed into textual representations. This involves converting documents into machine-readable formats using OCR systems to extract features such as text, bounding boxes, and visual elements Mieskes and Schmunk (2019); Smith (2007). Alternatively, a formatting language such as Markdown can be employed to represent the document’s layout, allowing the LLM to understand the structural context of the text better. The impact of OCR quality on IE performance has been documented Bhadauria et al. (2024), and structured formats tend to yield better results Bai et al. (2024). To process larger documents efficiently, they are often divided into smaller, manageable chunks based on page boundaries, sections, or semantic units Liu et al. (2024). Markdown as an input format compared to raw OCR outputs remains underexplored, representing a potential research gap in IE system development.
Model Engagement. Once preprocessed, the document is fed to an LLM for IE. Ensuring alignment between the extracted text and layout information is crucial for accurate representation Xu et al. (2020a); Appalaraju et al. (2021). Prompt-driven extraction leverages general-purpose models, using tailored prompts to guide the extraction process Brown et al. (2020); Radford et al. (2021); Zhou et al. (2022). As such, the model needs to be instructed to extract information, and usually, the schema is provided. This step can involve more advanced IT and ICL techniques (few-shot, CoT). The influence of prompting techniques in interaction with various stages of the IE pipeline to enhance performance and robustness remains a research gap that requires further investigation.
Output Refinement. After inference, the extracted information undergoes post-processing to ensure accuracy and conformance to Schema S. This step involves refining and validating the outputs generated by the LLM through tasks such as mapping extracted entities E to their original document positions, merging overlapping or fragmented predictions, and resolving ambiguities in the results Xu et al. (2020b). Refining entity extraction through post-processing has been explored in various studies Wang et al. (2022b); Tamayo et al. (2022). Rule-based entity alignment has notable accuracy improvements Luo et al. (2024). There exists no analysis of post-processing techniques tailored to LRDs in conjunction with LLMs.
3 Test-Suite for IE from Layout-rich Documents with LLMs
We implement a comprehensive test suite to assess IE tasks from LRDs. The pipeline, depicted in Figure 2, systematically transforms raw input into structured output across multiple stages for evaluating the efficacy of design decisions.
Figure 2: Overview of the test suite for extracting information from layout-rich documents using LLMs in six stages. The process begins with OCR-based text extraction and Markdown conversion with LLM assistance, followed by chunking to manage token limits, experimenting with different chunk sizes. Each chunk is processed to generate a prompt, exploring Few-shot and CoT prompting with varying numbers of examples. These examples consist of a document example and extracted key-value pairs in a structured format, along with a new document and the required task for the LLM. LLMs generate structured JSON outputs, which are then decoded and reconciled. Post-processing involves experimenting data cleaning and entity mapping. Finally, evaluation is conducted using two different evaluation techniques.
3.1 Data Structuring
We convert raw documents into machine-readable formats by extracting text and layout features. The process involves two conversions: 1) Extracting textual content and layout information using OCR data, and 2) creating a markdown representation of PDFs.
Chunking. We employ three chunk sizes: (1) max: 4096 tokens, (2) medium: 2048 tokens, and (3) small: 1024 tokens. Documents are segmented into $N$ chunks based on document length and chunk size, preparing them for the prompting phase. Each chunk is formed by accumulating whole words until the token limit. This ensures intact word boundaries and maintaining a sequential, non-overlapping structure. Layout information is preserved by associating each text segment with normalized and quantized spatial coordinates, retaining the structural context of the document.
3.2 Model Engagement
Model engagement consists of constructing input to the LLM comprising at least three components: (1) a task instruction outlining the IE task, (2) the target schema $S$, and (3) the document chunk. We adhere to best practices from NLP for prompt structure and IE task instruction. The schema $S$ is implemented as a dictionary of key-value pairs, where values specify the format of the corresponding attribute using regex expressions in Listing 1.
[…]
Listing 1: Schema Definition
We also implement two ICL strategies: (1) Few-Shot and (2) Chain-of-Thought (CoT) (see Appendix A.1 for more details). For $N$ prompts, the LLM performs N completions, with one completion per prompt. The outputs are collected and stored as raw predictions, ready for Output Refinement.
3.3 Output Refinement
We refine the raw outputs from the LLM to ensure alignment with the target schema, addressing challenges related to prediction variability, schema definition differences, and data formatting inconsistencies (e.g., varying date formats). We implement three techniques inspired by related work in data integration Dong and Srivastava (2013): Decoding, Schema Mapping, and Data Cleaning. This process results in three sets of predictions: initial predictions, mapped predictions, and cleaned predictions.
Decoding. The decoding step parses each LLM completion as a JSON object, discarding any that fail to parse. The process then consolidates predictions for each document by reconciling outputs generated across individual pages and chunks. With $N$ completions for $N$ prompts, corresponding to $N$ chunks, the model generates multiple predictions for a single document. Reconciliation ensures a unified document-level output by deduplicating nested predictions and aggregating unique values. If multiple unique values exist for a single entity, they are stored together to preserve variability. The outcome of this step is referred to as the initial predictions.
Schema Mapping. LLMs are expected to return only keys $\{a_{1},a_{2},...,a_{k}\}$ specified in the target schema $S$. However, they may occasionally fail to return the keys as expected. E.g., ‘file date” is returned instead of “file_date”. Such LLM “overcorrection” can hinder strict schema conformance. As a countermeasure, we implement a post-processing step that maps the predicted keys to align with the target schema. Our mapping step integrates multiple weak-supervision signals, such as exact matching, partial matching, and synonym-based logic, inspired by the recent techniques for ontology alignment Fürst et al. (2023). The outcome of this step is the mapped predictions, where entity keys $\{a_{1},a_{2},...,a_{k}\}$ are standardized and fully aligned with the schema $S$.
Data Cleaning. A common issue concerns the format of the values $T_{k}$ of predicted key-value pairs $\{(a_{1},T_{1}),(a_{2},T_{2}),...,(a_{k},T_{k})\}$. We must standardize formats, such as dates and names, to align with the target schema $S$. One source of error is LLM hallucinations Ji et al. (2023); Huang et al. (2023); Xu et al. (2024b), while another problem is that information is often not aligned to a common format inside the source data. For instance, two documents might use two different formats for dates (“April 1992” vs “1992-04-01”). Additional issues include capitalization, redundant whitespace, or special characters. We utilize the regex-defined data types in our schema to automatically apply data cleaning functions. The outcome of this step is the cleaned predictions, representing the final fully normalized outputs.
Evaluation Techniques. Evaluating IE for LRDs requires comparing the extracted data against an annotated test dataset. We implement three metrics for this evaluation: exact match, substring match, and fuzzy match.
  • Exact Match searches for perfect alignment between predicted and ground truth values. A match is valid only when the values are identical. This strict approach is ideal for extracting specific, unambiguous entities like dates or numerical identifiers.
  • Substring Match checks whether ground truth values are fully contained within the predicted values as complete substrings, without being split or partially matched. It ensures all ground truth values appear in their entirety within predictions, making it effective for tasks such as extracting full names or addresses, where additional contextual details (e.g., titles like Mr. and Mrs.) may be included in the predictions without making the extraction incorrect.
  • Fuzzy Match uses similarity metrics for approximate matches. A match is valid if the highest similarity ratio exceeds a predefined threshold (default: 0.8). This method is well-suited for scenarios with minor variations caused by OCR errors or formatting discrepancies.
4 Experimental Evaluation
4.1 Experimental Design
Methodology.
The goal of our experimental setup is to study how different parameters in the pipeline, shown in Figure 2, affect the overall performance of IE from LRDs using LLMs. To investigate the design dimensions, we start with a baseline configuration and and systematically alter factors at a single dimension at a time, following a one-factor-at-a-time (OFAT) methodology. This approach allows us to isolate and understand the impact of each parameter change on the IE performance.
Our intuition is that aggregating the knowledge gained for each dimension independently, we can achieve a deeper understanding of the design space and possibly identify an effective overall configuration for IE, without the need for a comprehensive factorial exploration. However, we also validate the findings by comparing the results of the OFAT method with those obtained from a brute-force approach, which is based on conducting 432 experiments.
Dataset and LLMs.
We utilize the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023b), which includes two benchmarks. Each benchmark includes training samples of 10, 50, 100, and 200 documents with high-quality OCR for assessing data efficiency, as well as generalization tasks: Single Template Learning (STL), Unseen Template Learning (UTL), and Mixed Template Learning (MTL). For further details on how this dataset is tailored for our experiments and diverse models, please refer to Appendix A.2.
We evaluate GPT-3.5, GPT-4o, and LLaMA3-70B for text-only structured data extraction from LRDs. Additionally, we compare their results with GPT-4 Vision and LayoutLMv3 to assess the performance gap between multimodal LLMs and domain-specific, fine-tuned models, respectively.
The Baseline Configuration.
The baseline configuration is outlined in Table 1,where the configuration is selected based on best practices such as in Perot et al. (2024) for the following reasons: (1) OCR reflects real-world scenarios for digitized LRDs. (2) Medium chunk size balances efficiency and context preservation, addressing token limits in LLMs. (3) Few-shot prompting combines pre-trained knowledge with minimal task-specific guidance. (4) Using zero examples provides a clear benchmark for assessing the model’s raw performance. (5) Initial predictions are retained to evaluate models’ raw output without modifications, ensuring a direct assessment of their capabilities. (6) Finally, exact match provides a stringent measure of correctness, offering a reliable baseline for comparison across configurations.
Table 1: Overall configuration parameters. Baseline configuration is highlighted with light blue.
Parameter Values
Input Type OCR, Markdown
Chunk Size Category Small, Medium, Max
Prompt Type Few-Shot, CoT
Example Number 0, 1, 3, 5
Post-processing Strategy Initial, Mapped, Cleaned
Evaluation Technique Exact, Substring, Fuzzy
4.2 The Input Dimension
We substitute OCR input with Markdown and as outcomes in both STL and UTL scenarios. The differences in performance between OCR and Markdown are model- and context-dependent, exhibiting no consistent trend favoring one input type over the other, as shown in Table 2.
Table 2: Performance results for different LLMs across STL and UTL levels with different input types. Baseline configuration in light blue.
Models Level Exact Match (F1)
OCR Markdown
GPT-3.5 STL 0.650 0.647 (-0.003)
UTL 0.645 0.657 (+0.012)
GPT-4o STL 0.670 0.633 (-0.037)
UTL 0.659 0.633 (-0.026)
LLaMA3 STL 0.640 0.657 (+0.017)
UTL 0.640 0.662 (+0.022)
Avg (±stdev.) 0.650 (±0.011) 0.648 (±0.012)
OCR input serves as a stable baseline for IE tasks, delivering consistent performance across models. GPT-4o has noticeable performance drops with Markdown input, indicating its reliance on OCR for optimal results. In contrast, Markdown marginally improves performance for LLaMA3-70B at both STL and UTL scenarios, suggesting its potential benefits from the additional structure or semantic cues. GPT-3.5 demonstrates robustness to changes in input type, with only slight fluctuations in performance. On average, OCR marginally outperforms Markdown (0.650 vs. 0.648), but the differences are minor, with standard deviations indicating similar stability.
4.3 The Chunk Dimension
To evaluate the impact of chunk size, we varied it from medium to max and small while keeping all other parameters constant. Table 3 demonstrates how chunk size affects performance across STL and UTL levels.
Table 3: Performance results for different LLMs across STL and UTL levels with different chunk size categories. Baseline configuration in light blue.
Models Level Exact Match (F1)
Small($\leq 1024$) Medium($\leq 2048$) Max($\leq 4096$)
GPT-3.5 STL 0.562(-0.088) 0.650 0.645(-0.005)
UTL 0.561(-0.084) 0.645 0.644(-0.001)
GPT-4o STL 0.602(-0.068) 0.670 0.674(+0.004)
UTL 0.600(-0.059) 0.659 0.657(-0.002)
LLaMA3 STL 0.615(-0.025) 0.640 0.647(+0.007)
UTL 0.608(-0.032) 0.640 0.644(+0.004)
Avg(±stdev.) 0.591(±0.023) 0.650(±0.011) 0.651(±0.011)
Medium and max chunk sizes provide the most consistent and stable results across models, with an average F1 score of 0.650 (±0.011) and 0.651 (±0.011), respectively. Due to insufficient context, small chunk sizes result in significant performance drops, particularly for GPT-3.5 and GPT-4o. These findings suggest that max chunk size is optimal, but medium can be a good option for LLMs with limited context lengths..
4.4 The Prompt Dimension
Table 4 presents the impact of prompt type and the number of examples on model performance at STL and UTL levels. Surprisingly, in-context demonstrations do not enhance performance for either few-shot or CoT experiments. For both experiments, the setting with zero examples achieves the highest average performance: few-shot 0.650 (±0.011) and CoT 0.649 (±0.008). Performance consistently declines as the number of examples increase, likely due to noise that impairs generalization. Overall, there is no significant difference between few-shot and CoT.
Table 4: Different LLMs across STL and UTL levels with different prompt types and example numbers. Baseline Configuration is highlighted in light blue.
Models Level Exact Match (F1)
0 1 3 5
few-shot
GPT-3.5 STL 0.650 0.586(-0.064) 0.593(-0.057) 0.548(-0.102)
UTL 0.645 0.566(-0.079) 0.564(-0.081) 0.541(-0.104)
GPT-4o STL 0.670 0.608(-0.062) 0.602(-0.068) 0.595(-0.075)
UTL 0.659 0.597(-0.062) 0.607(-0.052) 0.601(-0.058)
LLaMA3 STL 0.640 0.599(-0.041) 0.606(-0.034) 0.603(-0.037)
UTL 0.640 0.582(-0.058) 0.601(-0.039) 0.597(-0.043)
Avg(±stdev.) 0.650(±0.011) 0.589(±0.014) 0.595(±0.016) 0.580(±0.028)
CoT
GPT-3.5 STL 0.653(+0.003) 0.602(-0.048) 0.544(-0.106) 0.533(-0.117)
UTL 0.650(+0.005) 0.575(-0.007) 0.548(-0.097) 0.516(-0.129)
GPT-4o STL 0.655(-0.015) 0.615(-0.055) 0.612(-0.058) 0.605(-0.065)
UTL 0.659(0) 0.614(-0.045) 0.611(-0.048) 0.607(-0.052)
LLaMA3 STL 0.635(-0.005) 0.603(-0.037) 0.613(-0.027) 0.610(-0.003)
UTL 0.644(+0.004) 0.586(-0.054) 0.601(-0.039) 0.598(-0.042)
Avg(±stdev.) 0.649(±0.008) 0.599(±0.015) 0.588(±0.032) 0.578(±0.042)
4.5 Output Refinement
We examine two output refinement strategies, Schema Mapping and Data Cleaning (see Sec. 3.3), to evaluate their impact shown in Table 5.
Table 5: Different LLMs across STL and UTL levels with different post-processing strategies. Baseline configuration is highlighted in light blue.
Models Level Exact Match (F1)
Initial Pred. Mapped Pred. Cleaned Pred.
GPT-3.5 STL 0.650 0.650(0) 0.737(+0.087)
UTL 0.645 0.645(0) 0.733(+0.088)
GPT-4o STL 0.670 0.670(0) 0.749(+0.079)
UTL 0.659 0.659(0) 0.741(+0.082)
LLaMA3 STL 0.640 0.640(0) 0.724(+0.084)
UTL 0.640 0.640(0) 0.725(+0.085)
Avg(±stdev.) 0.650(±0.011) 0.650(±0.011) 0.734(±0.009)
Schema Mapping
involves mapping the predicted schema keys to the target schema keys. Our results show no change in F1 scores compared to the initial predictions. This suggests that the models already effectively return the correct attributes, making the mapping step unnecessary.
Data Cleaning
uses the defined data types to perform automatic value cleaning, consistently achieving the highest F1 scores across all models. This underscores the need for post-processing steps for IE with LLMs to align the extracted data with the target format to handle LLM hallucinations and inconsistent source data formats (see Sec. 3.3).
4.6 Evaluation Techniques
We explore three evaluation techniques to assess their impact on model performance (see Sec. 3.3). On average, Fuzzy Match achieved the highest F1 score (0.733), outperforming Substring Match (0.676) and Exact Match, as shown in Table 6. We provide a detailed error analysis of fuzzy and substring match accuracy in Appendix A.3, showing that they provide a near-perfect precision when manually checked for semantic equivalence with precision scores of 0.98 and 1.00, respectively. This shows Fuzzy Match’s ability to balance flexibility and precision.
Table 6: Different LLMs across STL and UTL levels with different evaluation techniques. Baseline configuration in light blue.
Models Level Exact Match Substring Match Fuzzy Match
(F1) (F1) (F1)
GPT-3.5 STL 0.650 0.683(+0.033) 0.730(+0.080)
UTL 0.645 0.682(+0.037) 0.726(+0.081)
GPT-4o STL 0.670 0.690(+0.020) 0.750(+0.080)
UTL 0.659 0.678(+0.019) 0.744(+0.085)
LLaMA3 STL 0.640 0.661(+0.021) 0.727(+0.087)
UTL 0.640 0.662(+0.022) 0.723(+0.083)
Avg(±stdev.) 0.650(±0.011) 0.676(±0.011) 0.733(±0.010)
4.7 Putting it All Together
In the preceding sections, we investigated the influence of various parameters on model performance along the IE extraction pipeline, analyzing one factor at a time. Drawing from the underlying 12 experiments, we identified the optimal parameter for each step and each model based on the experimental outcomes (Table 7). In addition, we conducted an exhaustive full factorial exploration with 432 configurations ($2*3*2*4*3*3=432$, see Table 1) to find the best parametrization per LLM (Table 8). Lastly, we find the worst configuration on a per LLM and per model basis (Table 9). The performance of these different configurations is depicted in Figure 3. We gain several insights:
  • OFAT approximates well the Brute-Force configuration with a fraction ($\sim 2.8\%$) of the required computation. We see in Table 7 and Table 8 that they match except for 4 parameter choices (prompt and example No. parameters for GPT-4o and LlaMa3). Likewise, their F1 scores are close to each other (Figure 3), with OFAT achieving 0.791 and Brute-Force achieving 0.801 overall.
  • Adapting the IE pipeline to the LLM is necessary to achieve a competitive performance. Overall, the OFAT configuration improves from the baseline F1 of 0.650 to 0.791, a $22\%$ improvement. For Brute-Force, the improvement is $23\%$. In comparison, the worst configuration achieves only 0.45 on average, almost half of the best configuration.
  • Besides the need for pipeline customization, common patterns exist across LLMs. First, output refinement and evaluation techniques boost performance across all models. Second, there is a tendency to larger context sizes. Lastly, larger models (GPT-4o, LLaMa3) benefit more from examples, while the CoT pattern generally aids IE.
Table 7: OFAT configurations on a per-model basis and corresponding performance results.
Parameter GPT-3.5 GPT-4o LLaMA3
Input Type Markdown OCR Markdown
Chunk Size Medium Max Max
Prompt CoT Few-Shot Few-Shot
Example No. 0 0 0
Output Refin. Cleaned Cleaned Cleaned
Evaluation Fuzzy Fuzzy Fuzzy
Table 8: Brute-Force configurations on a per-model basis and corresponding performance results.
Parameter GPT-3.5 GPT-4o LLaMA3
Input Type Markdown OCR Markdown
Chunk Size Medium Max Max
Prompt CoT CoT CoT
Example No. 0 5 5
Output Refin. Cleaned Cleaned Cleaned
Evaluation Fuzzy Fuzzy Fuzzy
Table 9: Worst performance configurations on a per-model basis and corresponding performance results.
Parameter GPT-3.5 GPT-4o LLaMA3
Input Type OCR Markdown OCR
Chunk Size Small Small Small
Prompt Few-shot Few-shot Few-shot
Example No. 5 5 5
Output Refin. Initial Initial Initial
Evaluation Exact Exact Exact
Figure 3: F1 scores of different LLMs across three configurations (Baseline, OFAT, and Brute-Force). Each bar represents the mean F1 score of a model for the corresponding configuration.
Finally, we compare our purely text-based approach to IE to (1) GPT4-vision, a multimodal LLM, and (2) LayoutLMv3, a leading layout-aware model that we fine-tune to our dataset (Table 10). Despite fine-tuning, LayoutLMv3 does not perform as well as our evaluated LLMs, even on documents that contain the same structure as the fine-tuning data (STL). We also observe that LLMs remain stable across template types. The best-performing model is GPT4-vision, a multi-modal LLM, where we directly provide an image of the PDF for IE. The model benefits from less context loss due to chunking and its ability to jointly use textual and visual features. However, when considering token usage and cost, GPT4-vision requires $\sim 2$ times the tokens compared to text-only approaches and $>10$ times the cost under the current OpenAI pricing scheme (Nov. 2024). Therefore, text-only approaches constitute a good trade-off between performance and cost for practical applications.
Table 10: Cost (per experiment) and performance of LLMs across STL and UTL levels for the OFAT configuration. The tokens for LayoutLMv3 include the fine-tuning. We run LauoutLMv3 and LlaMa3 locally, without API costs.
Models Tokens API Cost STL (F1) UTL (F1)
GPT-3.5 322K \$0.18 0.778 0.754
GPT-4o 278K \$0.40 0.797 0.790
LLaMA3 239K \$0 0.843 0.820
GPT-4-vision 585K \$4.76 0.902 0.897
LayoutLMv3 176.1M \$0 0.603 0.194
5 Related Work
IE using LLMs.
Transformer-based models like BERT Devlin et al. (2019), GPT Brown et al. (2020), and RoBERTa Liu et al. (2019b) have advanced NLP with self-attention and large-scale pretraining but struggle with layout-rich documents (LRDs). To address this, layout-aware and multimodal models have emerged. LayoutLM Xu et al. (2020b) integrates spatial features, with later versions Xu et al. (2021, 2022) and models like GPT-4V OpenAI (2023) and Gemini Pro Anil et al. (2023) enhancing document understanding through multimodal learning. LayoutLLM Fujitake (2024) and structural-aware approaches Lee et al. (2022) further improve extraction accuracy, while end-to-end models like Donut Kim et al. (2022) bypass OCR for direct document image processing. Additionally, ChatUIE Xu et al. (2024a) adopts a chat-based approach for flexible IE, while ReLayout Jiang et al. (2025) enhances document understanding through layout-aware pretraining, advancing LLMs for real-world use.
Strategies for IE from LRDs.
Graph-based models like GCNs Liu et al. (2019a) and AliGATr Nourbakhsh et al. (2024) enhance relation extraction by capturing textual-visual relationships. Reading order is critical; Token Path Prediction (TPP) Zhang et al. (2023) resolves OCR layout ambiguities, while global tagging Shaojie et al. (2023) mitigates text ordering issues for better extraction. For structured data, TabbyPDF Jain et al. (2020) targets table extraction, DocExtractor Zhong et al. (2020) processes forms, and LMDX Perot et al. (2024) unifies OCR, preprocessing, and postprocessing for document IE. Additionally, XFORMPARSER Cheng et al. (2025) offers a simple yet effective multimodal and multilingual approach for parsing semi-structured forms.
Preprocessing, Chunking, Prompting, Postprocessing, and Evaluation Techniques.
Chain-of-Thought (CoT) prompting Wei et al. (2022) enhances reasoning in complex LRD extraction tasks, while diverse prompt-response datasets Zmigrod et al. (2024a) improve LLM robustness. Instruction-finetuned LLMs have also demonstrated effectiveness in domain-specific applications, such as clinical and biomedical tasks, where zero-shot and few-shot learning enable adaptive extraction without extensive fine-tuning Labrak et al. (2024). Postprocessing techniques, including text normalization, entity resolution Hwang et al. (2021), and majority voting Wang et al. (2022a), refine extracted data by correcting OCR and extraction errors for greater accuracy.
Despite advancements, current studies often focus on isolated components rather than evaluating full IE pipelines. Key stages such as OCR, chunking, and postprocessing are frequently assessed independently, leading to an incomplete understanding of their interplay. Our research bridges this gap by systematically evaluating strategies across all pipeline stages to identify optimal configurations for leveraging pre-trained models.
6 Conclusion
In this paper we introduce the design space for IE from LRDs using LLMs, with challenges of data structuring, model engagement, and output refinement. We consider the design choices within those challenges. We examine the effects of these choices on the overall performance through OFAT and full-factorial exploration of the parameters. Finally, we show that a well-designed general purpose LLM can provide a cost-effective alternative to the the specialized models for IE from LRDs.
Limitations
For the experimental evaluation in the design space, we focused on the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023b), which has ground-truth annotations that make it easy for result analysis. We believe incrementally adding new datasets focusing on layout-rich documents with ground truth and making a combined benchmark would help improve the design space itself as we can test different application-specific scenarios.
The evaluation metrics are planned to be extended with metrics that are popularly used in the context of LLMs such as ROUGE Lin (2004) and BLEU Papineni et al. (2002) scores and their adaptations Yang et al. (2018). A fair comparison using cost-related metrics is mostly harder to compute due to their changing depending on the scenario. For instance, in some scenarios where we listed cost as 0 due to having a free-to-use service, it actually comes with a cost. Similarly, the cost of computation can be considered from different perspectives such as the energy usage of a model.
The test suite is currently limited to the steps that we listed in Fig. 2, whereas one could imagine additional steps or factors which affect the performances and may even deliver more satisfactory outcomes. We would like to incorporate additional steps that we learn from the community and incrementally enlarge the design space and extend the testing capabilities for IE from LRDs.
Last, we have only evaluated three LLMs, one multi-modal LLM and one layout-aware, fine-tuned model (LayoutLMv3). In the future, we want to extend our evaluation to further LLMs and fine-tuned, layout-aware models (e.g., ). Despite that, our results already show a clear trend in which specialized, fine-tuned models struggle to keep up with progress in general-purpose LLMs. E.g., LayoutLMv3, only reaches an F1 score of 0.681 with fine-tuning and on the same document types (STL), substantially less than Llama3, the best text-based LLM with 0.832 and GPT-4-vision, our tested multi-modal LLM with 0.890. This underlines the importance of our work, in which choices in data structuring, model engagement, and output refinement become critical in achieving the best performance.
Ethical Considerations
Large Language Models (LLMs) can contain biases that can have a negative impact on marginalized groups Gallegos et al. (2024). For the task of information extraction, this could have the impact that uncommon names for people and places are auto-corrected by the LLM to their more common form. In our experiments, we have encountered some instances of this and plan to investigate this further.
Appendix A Appendix
A.1 Prompt Generation Details
Document representation in the example section of the prompt is generated by the LLM, condensing the original OCR data while retaining information relevant to the target schema to reduce token usage and cost. This version integrates only text features. However, in the new example section, both text and layout features are preserved without modification, embedding spatial structure into text sequences to enhance model understanding. This is the only part that differs between the example and new document sections of the prompt. Task descriptions provide clear extraction guidelines, specifying what information to retrieve, while schema representation defines the expected JSON format to ensure consistency in extracted data. Additionally, CoT prompting includes a reasoning component, guiding the model through logical steps to improve accuracy on complex tasks (see Figures 45).
Prompt Template: Few-shot ### Examples ### (<Document>) {DOCUMENT_REPRESENTATION} (</Document>) (<Task>) {TASK_DESCRIPTION} {SCHEMA_REPRESENTATION} (</Task>) (<Extraction>) {EXTRACTION} (</Extraction>) ### New Documents ### (<Document>) {DOCUMENT_REPRESENTATION} (</Document>) (<Task>) {TASK_DESCRIPTION} {SCHEMA_REPRESENTATION} (</Task>) (<Extraction>)
Figure 4: Few-Shot Prompt Structure with 1-shot Example
Prompt Template: CoT ### Examples ### (<Document>) {DOCUMENT_REPRESENTATION} (</Document>) (<Task>) {TASK_DESCRIPTION} {SCHEMA_REPRESENTATION} (</Task>) (<Reasoning>) {REASONING} (</Reasoning>) (<Extraction>) {EXTRACTION} (</Extraction>) ### New Documents ### (<Document>) {DOCUMENT_REPRESENTATION} (</Document>) (<Task>) {TASK_DESCRIPTION} {SCHEMA_REPRESENTATION} (</Task>) (<Reasoning>) {REASONING} (</Reasoning>) (<Extraction>)
Figure 5: Chain of Thought Prompt Structure with 1-shot Example
A.2 Tailoring Dataset for our Test-Suite
We used the Registration Form with six entity types for this project, as shown in Figure 6.
"file_date": "",
"foreign_principle_name": "",
"registrant_name": "",
"registration_num": "",
"signer_name": "",
"signer_title": ""
Figure 6: VRDU Registration Form Entities.
The VRDU dataset includes predefined few-shot splits consisting of training, testing, and validation sets. These splits contain 10, 50, 100, and 200 training samples, each with three subsets, as shown in Figure 7. The dataset also includes different levels (Lv1: Single, Lv2: Mixed, and Lv3: Unseen Type) and various template types (Amendment, Dissemination, and Short-Form).
Figure 7: Representation of few_shot_splits from the VRDU dataset.
For each template-level combination, we selected the first JSON file ending in 0 with 10 training samples. Since this training data will be utilized for few-shot and Chain-of-Thought (CoT) prompting, only the first five documents were chosen from the training samples of the selected JSON files. Each level type (STL, UTL) includes template types (Amendment, Dissemination, Short-Form), each with 0, 1, 3, or 5 examples. In STL, these categories use the first document for one-example prompts, the first three for three-example prompts, and all five for five-example prompts. The same structure applies to UTL, with examples specific to its categories. This ensures consistency across template-level combinations while varying the number of examples in the prompt. Figure 8 shows an example of few-shot and CoT examples. This process was repeated for every level, template type, and example count. The example texts were generated using a Large Language Model (LLM), which was instructed to summarize the provided OCR text for the given document while ensuring the inclusion of target schemas and entities.
few_shot_examples = {
"STL": {
"Amendment": {
0: [],
1: [
"text": "This document is an amendment to the regis...",
"entities": {
"file_date": "1982-10-31",
"foreign_principle_name": "Japan Trade Center...",
"registrant_name": "PressAid Center",
"registration_num": "1833",
"signer_name": "Akira Tsutsumi",
"signer_title": "Director General"
3: [
5: [
"Dissemination": {...
Figure 8: VRDU Registration Form Entities.
For the test files, to ensure a fair comparison, we selected the first 40 common files from the chosen JSON files within each template type at each level. This means that for Lv1 Amendment and Lv3 Amendment test files, the first common 40 files were selected as test files, and the same strategy was applied for other template types as well. Due to the mixed nature of test files in Lv2, the mixed template type was excluded from this project.
For the models GPT-3.5, GPT-4o, and LLaMA3-70B, we created these few-shot and CoT examples. However, no few-shot examples were used for GPT-4 Vision; we used the same test dataset as for the other three models with basic instructions. For LayoutLMv3, we used the same training and test datasets as the other models but included the entire validation set (300 samples) instead of selecting only five. The training and test datasets remained unchanged across all models to ensure a fair evaluation and consistency.
A.3 Success of Evaluation Techniques
We manually reviewed the results of the baseline experiment from GPT-3.5, GPT-4o, and LLaMA3-70B to assess the success of the substring and fuzzy match metrics. The analysis focused on cases where the exact match score was 0 but substring/fuzzy was 1, highlighting predictions that failed strict matching but were successfully handled in other techniques. This created a dataset to test how well substring and fuzzy matching handle difficult cases. In total, we examined 91 key-value pairs for fuzzy match and 37 for substring match.
Figure 9: Comparison of error rates in Substring vs. Fuzzy Matching
Based on our intubation, we came up with seven different error categories see Figure 9. From left to right in Error Types 5 OCR error means errors occur because of OCR; for example, handwritten names’ or signatures are sometimes extracted as different names incorrectly (e.g Jim Slattery handwriting extracted as Jim Slatters) GT error represents the case where GT information is not completely true sometimes its missing some words and sometimes it has additional words for example gt value is ’Om Saudi Arabia 1’ but it is Kingdom’ Saudi Arabia in the document. LLM Halucination represents where LLM come up a prediction that is not exist data in the OCR for example gt value is for file date ’1992-04-21’ but prediction is ’1992-04-24’ even though OCR data and document itself does not contain any number 24. Additional info represents that prediction is correct, but there is some additional info’ for example, while the signer’s name is ’Daniel Manatt” LLM predicts it as ’Daniel Manatt Todd’ where Todd exists in OCR and the document has a notary name as a handwritten signature. Wrong info means gt is exist in the OCR and the document but LLM choose another date as prediction for example while file date ’2016-10-08’ exist in OCR LLM return 2016-10-31 as file date which belongs registration date in the document.
Table 11: Performance results of substring and fuzzy evaluation techniques over exact match based on manually annotated data, considering categorized error types.
Data Evaluation Precision Recall F1
Points Techniques
37 Exact Match 0.000 0.000 0.000
Substring Match 1.000 1.000 1.000
91 Exact Match 0.000 0.000 0.000
Fuzzy Match 0.984 1.000 0.992
Human error is for the errors occurred physically by humans failures for example sometimes documents ar scanned partly/crooked by humans. Finally incomplete predictions represent prediction is include gt but not completely for example sometime gt is ’Japan External Trade Organization Tokyo, Japan’ but prediction is missing only last part which is Tokyo, Japan. Figure 9 represent all these error types and their rates in both Substring and Fuzzy match categories where exact match is 0 but they labeled as 1.
For our evaluation the success of Fuzzy and Substring we decided to label data points as 0 only in the case where wrong info is 1. For other error categories we accept them and label gt as 1 for those. Table 11 represent performance metrics for fuzzy and substring calculated with gt labels explained above with the data points (exact match 0 substring/fuzzy 1)